# Better Call SAL: Towards Learning to Segment Anything in Lidar

### v1

@@ Appendix E Qualitative Results
Class-agnostic segmentation.  In Fig.˜E.2, we provide qualitative results for class-agnostic segmentation on Waymo Open dataset. Colors encode identities (IDs) of individual instances. For reference, we provide images of corresponding camera views, even though these are not used during the inference. As can be seen, SAL learns to segment full point clouds, things and stuff classes, even though supervision is only partial. Interestingly, SAL segments well large structures (often classified as stuff classes), such as building, road, and sidewalk. In addition to canonical things classes, such as car, van, bus, and pedestrian, SAL also learns a variety of classes, that are not covered in class vocabularies of existing datasets of Lidar Panoptic Segmentation (SemanticKITTI [6] and nuScenes [18]). Examples of such classes, segmented in Fig.˜E.2, are parking meters, potted trees, rooftop ladder, water hydrant, post box, traffic cone and traffic barrier.
@@
In Fig.˜E.1, we contrast a baseline that simply lifts SAM [29] masks to Lidar (this baseline is evaluated in Tab.˜B.3, as well as Tab.˜4 in the main paper). As can be seen in Fig.˜E.1, left, Waymo Open provides $270^{\circ}$ coverage (four camera views), leaving a “blind spot” behind the vehicle. This is an inherent limitation of the baseline that requires camera views for the inference. By contrast, SAL (Fig.˜E.1, right) learns to distill such image-generated pseudo-labels to a full Lidar Panoptic Segmentation model. Therefore, it segments full $360^{\circ}$ point clouds and does not require camera views during the inference (only during the model training).
@@
For further insights, we zoom in on certain regions of point clouds. As can be seen in Fig.˜E.1, left, masks transferred from images to Lidar often lead to bleeding edges (e.g., red and blue pedestrians “bleed” to the pale blue wall), and thinner structures are often not segmented well due to non-perfect calibration and rolling shutter nature of the Lidar sensor. While such issues can be (partially) mitigated by postprocessing via density-based clustering (we refer to Tab.˜B.3 and Tab.˜4 in the main paper for quantitative analysis), we show in Fig.˜E.1, right, results obtained with SAL model, trained directly on SAM-transferred masks (without DBSCAN as postprocessing). Remarkably, the distilled model does not suffer from these artifacts.
@@
Figure E.1: Class-Agnostic Segmentation on Waymo Open [75]. We visually compare class-agnostic segmentation results. Colors encode object instance IDs. Left: baseline (SAM [31], unprojected to Lidar), and, right, SAL. As can be seen, the baseline that directly lifts SAM masks to Lidar is limited to $270^{\circ}$ field of view, which overlaps with the camera ring. By contrast, SAL segments the full Lidar point cloud and is not limited by the camera coverage. Zoomed-in regions show that the baseline is sensitive to edge bleeding (e.g., see pedestrian and traffic sign masks, partially projected to the blue wall). SAL, by contrast, distills noisy SAM masks into crisp segmentation masks.
@@
Figure E.2: Class-agnostic segmentation on Waymo Open [75] from first-person perspective. We visually outline the Lidar point cloud, where points are colored according to estimated instance IDs, estimated by SAL. We show corresponding camera views (not used for inference) for reference. As can be seen, SAL accurately segments a large variety of objects, including parking meters, potted trees (pots as well as trees), rooftop ladder, water hydrant, post box, traffic cone, traffic barrier, and more. Canonical objects, such as car, van, bus, and pedestrian are segmented as well. This class-agnostic segmentation is a basis for zero-shot classification.
Zero-shot prompting.  In Fig.˜E.9, we highlight the capability of SAL for prompting specific semantic classes. On the left side, we visualize class-agnostic segmentations of point clouds, and on the right, we highlight prompts and highlight segmented regions. We show two things classes (tram and trash bin) and two stuff lasses (store front and curb). A basis for such zero-shot prompting is our class-agnostic segmentation model, which segments input point clouds into a set of segmented objects. SAL model predicts for each segmented object CLIP feature token that we use for zero-shot prompting (for details, we refer to the main paper, Sec.˜3).
Zero-shot semantic segmentation.  To provide further insight, we illustrate semantic ground truth labels, SAL pseudo-labels, that we use to train the SAL model, and SAL model outputs for SemanticKITTI [6] (Fig.˜E.6), nuScenes [18] (Fig.˜E.7) and Waymo Open [75] datasets.
@@
Figure E.3: SAM mask and CLIP vocabulary predictions - SemanticKITTI. We show the front camera view of the first scan of the $0$, 01, 02, and 08 sequences. We obtain masks with SAM [31] and compute per-mask CLIP [67] image tokens with MaskCLIP [16]. To visualize classes, we prompt generated tokens with SemanticKITTI [6] class vocabulary. These classes are not transferred to Lidar. Our pseudo-labels contain masks and image tokens, no explicit class labels. Instances of the same class are indicated in different tones of the same color.
@@
Figure E.4: SAM mask and CLIP vocabulary predictions - nuScenes. We show all six camera views of the first scan of the 0001 sequence. We obtain masks with SAM [31] and compute per-mask CLIP image tokens with MaskCLIP [16]. To visualize classes, we prompt generated tokens with nuScenes [18] class vocabulary. These classes are not transferred to Lidar. Our pseudo-labels contain masks and image tokens, no explicit class labels. Instances of the same class are indicated in different tones of the same color.
@@
Figure E.5: SAM mask and CLIP vocabulary predictions - Waymo. We show all five camera views of the first frame of sequence emph008 of the test set. We obtain masks with SAM [31] and compute per-mask CLIP image tokens with MaskCLIP [16]. Sine WAYMO has no panoptic ground truth, we visualize classes by prompting generated tokens with the nuScenes [18] class vocabulary. These classes are not transferred to Lidar. Our pseudo-labels contain masks and image tokens, no explicit class labels. Instances of the same class are indicated in different tones of the same color.
Table E.1: Dataset vocabulary text prompts. To circumvent ambiguous or uninformative class names, we prompt each class with a set of possible text prompts. Furthermore, we follow [67] and wrap each class prompt in full-sentence templates.
@@
Figure E.6: Qualitative results for SemanticKITTI. We visualize ground truth (GT), pseudo labels, and our model output for several scans of validation sequence 08 of SemanticKITTI [6]. While GT and our output display semantics, the class-agnostic pseudo labels show instances.
@@
Figure E.7: Qualitative results for nuScenes. We visualize ground truth (GT), pseudo labels, and our model output for the first scan of validation sequences 0003, 0013, 0015, 0017, 0035, and 0038 of nuScenes [18]. While GT and our output display semantics, the class-agnostic pseudo labels show instances.
@@
Figure E.8: Qualitative results for Waymo. We visualize pseudo labels and our model output for the first scan of test sequences 008, 024, and 032 of Waymo [75]. Waymo does not provide panoptic ground truth labels. While our output displays semantics, the class-agnostic pseudo labels show instances.
@@
Figure E.9: Zero-shot per-class prompting on Waymo Open [75]. SAL predicts a set of object instances (left), along with their objectness scores and distilled CLIP [67] features. We can use text prompts and query these instances for specific classes specified as prompts. On the right, we highlight several such examples that are outside of class-vocabularies of SemanticKITTI, nuScenes, and Waymo Open datasets. As can be seen on the left, a basis for such zero-shot prompting is accurate and, importantly, diverse class-agnostic segmentation.

### v2

@@ Appendix 0.E Qualitative Results
Class-agnostic segmentation.  In Fig.˜0.E.2, we provide qualitative results for class-agnostic segmentation on Waymo Open dataset. Colors encode identities (IDs) of individual instances. For reference, we provide images of corresponding camera views, even though these are not used during the inference. As can be seen, SAL learns to segment full point clouds, things and stuff classes, even though supervision is only partial. Interestingly, SAL segments well large structures (often classified as stuff classes), such as building, road, and sidewalk. In addition to canonical things classes, such as car, van, bus, and pedestrian, SAL also learns a variety of classes, that are not covered in class vocabularies of existing datasets of Lidar Panoptic Segmentation (SemanticKITTI [6] and nuScenes [18]). Examples of such classes, segmented in Fig.˜0.E.2, are parking meters, potted trees, rooftop ladder, water hydrant, post box, traffic cone and traffic barrier.
@@
In Fig.˜0.E.1, we contrast a baseline that simply lifts SAM [29] masks to Lidar (this baseline is evaluated in Tab.˜0.B.3, as well as Tab.˜4 in the main paper). As can be seen in Fig.˜0.E.1, left, Waymo Open provides $270^{\circ}$ coverage (four camera views), leaving a “blind spot” behind the vehicle. This is an inherent limitation of the baseline that requires camera views for the inference. By contrast, SAL (Fig.˜0.E.1, right) learns to distill such image-generated pseudo-labels to a full Lidar Panoptic Segmentation model. Therefore, it segments full $360^{\circ}$ point clouds and does not require camera views during the inference (only during the model training).
@@
For further insights, we zoom in on certain regions of point clouds. As can be seen in Fig.˜0.E.1, left, masks transferred from images to Lidar often lead to bleeding edges (e.g., red and blue pedestrians “bleed” to the pale blue wall), and thinner structures are often not segmented well due to non-perfect calibration and rolling shutter nature of the Lidar sensor. While such issues can be (partially) mitigated by postprocessing via density-based clustering (we refer to Tab.˜0.B.3 and Tab.˜4 in the main paper for quantitative analysis), we show in Fig.˜0.E.1, right, results obtained with SAL model, trained directly on SAM-transferred masks (without DBSCAN as postprocessing). Remarkably, the distilled model does not suffer from these artifacts.
@@
Figure 0.E.1: Class-Agnostic Segmentation on Waymo Open [75]. We visually compare class-agnostic segmentation results. Colors encode object instance IDs. Left: baseline (SAM [31], unprojected to Lidar), and, right, SAL. As can be seen, the baseline that directly lifts SAM masks to Lidar is limited to $270^{\circ}$ field of view, which overlaps with the camera ring. By contrast, SAL segments the full Lidar point cloud and is not limited by the camera coverage. Zoomed-in regions show that the baseline is sensitive to edge bleeding (e.g., see pedestrian and traffic sign masks, partially projected to the blue wall). SAL, by contrast, distills noisy SAM masks into crisp segmentation masks.
@@
Figure 0.E.2: Class-agnostic segmentation on Waymo Open [75] from first-person perspective. We visually outline the Lidar point cloud, where points are colored according to estimated instance IDs, estimated by SAL. We show corresponding camera views (not used for inference) for reference. As can be seen, SAL accurately segments a large variety of objects, including parking meters, potted trees (pots as well as trees), rooftop ladder, water hydrant, post box, traffic cone, traffic barrier, and more. Canonical objects, such as car, van, bus, and pedestrian are segmented as well. This class-agnostic segmentation is a basis for zero-shot classification.
Zero-shot prompting.  In Fig.˜0.E.9, we highlight the capability of SAL for prompting specific semantic classes. On the left side, we visualize class-agnostic segmentations of point clouds, and on the right, we highlight prompts and highlight segmented regions. We show two things classes (tram and trash bin) and two stuff lasses (store front and curb). A basis for such zero-shot prompting is our class-agnostic segmentation model, which segments input point clouds into a set of segmented objects. SAL model predicts for each segmented object CLIP feature token that we use for zero-shot prompting (for details, we refer to Sec.˜3 of the main paper).
Zero-shot semantic segmentation.  To provide further insight, we illustrate semantic ground truth labels, SAL pseudo-labels, that we use to train the SAL model, and SAL model outputs for SemanticKITTI [6] (Fig.˜0.E.6), nuScenes [18] (Fig.˜0.E.7) and Waymo Open [75] datasets.
@@
Figure 0.E.3: SAM mask and CLIP vocabulary predictions - SemanticKITTI. We show the front camera view of the first scan of the $0$, 01, 02, and 08 sequences. We obtain masks with SAM [31] and compute per-mask CLIP [67] image tokens with MaskCLIP [16]. To visualize classes, we prompt generated tokens with SemanticKITTI [6] class vocabulary. These classes are not transferred to Lidar. Our pseudo-labels contain masks and image tokens, no explicit class labels. Instances of the same class are indicated in different tones of the same color.
@@
Figure 0.E.4: SAM mask and CLIP vocabulary predictions - nuScenes. We show all six camera views of the first scan of the 0001 sequence. We obtain masks with SAM [31] and compute per-mask CLIP image tokens with MaskCLIP [16]. To visualize classes, we prompt generated tokens with nuScenes [18] class vocabulary. These classes are not transferred to Lidar. Our pseudo-labels contain masks and image tokens, no explicit class labels. Instances of the same class are indicated in different tones of the same color.
@@
Figure 0.E.5: SAM mask and CLIP vocabulary predictions - Waymo. We show all five camera views of the first frame of sequence emph008 of the test set. We obtain masks with SAM [31] and compute per-mask CLIP image tokens with MaskCLIP [16]. Sine WAYMO has no panoptic ground truth, we visualize classes by prompting generated tokens with the nuScenes [18] class vocabulary. These classes are not transferred to Lidar. Our pseudo-labels contain masks and image tokens, no explicit class labels. Instances of the same class are indicated in different tones of the same color.
Table 0.E.1: Dataset vocabulary text prompts. To circumvent ambiguous or uninformative class names, we prompt each class with a set of possible text prompts. Furthermore, we follow [67] and wrap each class prompt in full-sentence templates.
@@
Figure 0.E.6: Qualitative results for SemanticKITTI. We visualize ground truth (GT), pseudo labels, and our model output for several scans of validation sequence 08 of SemanticKITTI [6]. While GT and our output display semantics, the class-agnostic pseudo labels show instances.
@@
Figure 0.E.7: Qualitative results for nuScenes. We visualize ground truth (GT), pseudo labels, and our model output for the first scan of validation sequences 0003, 0013, 0015, 0017, 0035, and 0038 of nuScenes [18]. While GT and our output display semantics, the class-agnostic pseudo labels show instances.
@@
Figure 0.E.8: Qualitative results for Waymo. We visualize pseudo labels and our model output for the first scan of test sequences 008, 024, and 032 of Waymo [75]. Waymo does not provide panoptic ground truth labels. While our output displays semantics, the class-agnostic pseudo labels show instances.
@@
Figure 0.E.9: Zero-shot per-class prompting on Waymo Open [75]. SAL predicts a set of object instances (left), along with their objectness scores and distilled CLIP [67] features. We can use text prompts and query these instances for specific classes specified as prompts. On the right, we highlight several such examples that are outside of class-vocabularies of SemanticKITTI, nuScenes, and Waymo Open datasets. As can be seen on the left, a basis for such zero-shot prompting is accurate and, importantly, diverse class-agnostic segmentation.